OpenStudy (anonymous):

write as logarithm of a single quantity: 1/3[4log5(x+1)-3log5(3-x)+6log5x]

OpenStudy (nowhereman):

So how far do you get?

OpenStudy (anonymous):

not at all because its so confusing

OpenStudy (anonymous):

is this log base 5?

OpenStudy (anonymous):

actually it doesnt matter so lets just write log

OpenStudy (anonymous):

and ignore the 1/3 out front we will deal with it last

OpenStudy (anonymous):

yes its a base 5

OpenStudy (anonymous):

first use \[\log(a^n)=n\log(a)\] backwards. all the multipliers come inside the log as exponents

OpenStudy (anonymous):

\[4\log(x+1)=\log((x+1)^4)\] \[3\log(3-x)=\log((3-x)^3)\] \[6\log(x)=\log(x^6)\]

OpenStudy (anonymous):

giving \[\log((x+1)^4)-\log((3-x)^3)+\log(x^6)\]

OpenStudy (anonymous):

now we use \[\log(a)-log(b)=log(\frac{a}{b})\]

OpenStudy (anonymous):

what happened to the 1/3 outside the []

OpenStudy (anonymous):

i will deal with that last

OpenStudy (anonymous):

not done yet!

OpenStudy (anonymous):

\[log((x+1)^4)-\log((3-x)^3)=\log(\frac{(x+1)^4}{(3-x)^3})\]

OpenStudy (anonymous):

now we use \[\log(a)+\log(b)=\log(ab)\]

OpenStudy (anonymous):

\[\log(\frac{(x+1)^4}{(3-x)^3})+\log(x^6)=\log(\frac{x^6(x+1)^4}{(3-x)^3})\]

OpenStudy (anonymous):

now for the 1/3 out front. take the cube root of all that stuff.

OpenStudy (anonymous):

the inner stuff i mean. the cube root of x^6 is x^2, and the cube root of (3-x)^3 is 3-x and so you get \[\log(\frac{x^2(x+1)^{\frac{4}{3}}}{3-x})\]

OpenStudy (anonymous):

wow that was a hard problem but thanks again:)